OpenStudy (anonymous):

given sec theta=3 . use trigonometric identities to find the exact value of : csc (90degrees-theta)

OpenStudy (amistre64):

sec = 3/1 = 1/cos cos = 1/3; c^2+s^2 = 1 sin = sqrt(1-1/9) = 2sqrt(2)/3 sin = 1/csc = 2sqrt(2)/3 csc = 3/2sqrt(2) the rest might be an application of the sin(a+b) stuff

OpenStudy (amistre64):

or, sin 90 = pi/2 csc (pi/2 - 3/2sqrt(2)) but that loks like a pain to me

OpenStudy (amistre64):

hmm, we want theta itself tho

OpenStudy (amistre64):

theta = arccos(1/3) ewww

myininaya (myininaya):